The perimeter of a rectangle is 13cm. If its length is four 3/5 cm, then find its breadth.
Answer by Cromlix(4381)

Perimeter = 2 x length + 2 x breadth
13cm = 2 x 4 3/5 + 2 x breadth
65/5 = 46/5 + 2 x breadth
65/5 - 46/5 = 2 x breadth
19/5 = 2 x breadth
Divide both sides by 2
19/10 = breadth.
Proof:
Perimeter = 2 x length + 2 x breadth
= 2 x 23/5 + 2 x 19/10
= 46/5 + 19/5
= 65/5 cm
= 13 cm.
Hope this helps.
:-)